OpenStudy (_excusemyfrench._):

What is the equation of the line that passes through (0,4) and (8,0)?

OpenStudy (_excusemyfrench._):

@AihberKhan

OpenStudy (caozeyuan):

well, the y intercept is obviously 4

OpenStudy (caozeyuan):

so we have y=kx+4, and since k=dy/dx where dy=y2-y1 and dx=x2-x1

OpenStudy (caozeyuan):

0-4=-4 and 8-0=8, so dy/dx=-1/2

OpenStudy (caozeyuan):

now you have it y=-1/2x+4

OpenStudy (_excusemyfrench._):

thank you!!!